# Diffusion Models for Tabular Data Imputation and Synthetic Data Generation

### v1

@@ Abstract.
Data imputation and data generation have important applications for many domains, like healthcare and finance, where incomplete or missing data can hinder accurate analysis and decision-making. Diffusion models have emerged as powerful generative models capable of capturing complex data distributions across various data modalities such as image, audio, and time series data. Recently, they have been also adapted to generate tabular data. In this paper, we propose a diffusion model for tabular data that introduces three key enhancements: (1) a conditioning attention mechanism, (2) an encoder-decoder transformer as the denoising network, and (3) dynamic masking. The conditioning attention mechanism is designed to improve the model’s ability to capture the relationship between the condition and synthetic data. The transformer layers help model interactions within the condition (encoder) or synthetic data (decoder), while dynamic masking enables our model to efficiently handle both missing data imputation and synthetic data generation tasks within a unified framework. We conduct a comprehensive evaluation by comparing the performance of diffusion models with transformer conditioning against state-of-the-art techniques, such as Variational Autoencoders, Generative Adversarial Networks and Diffusion Models, on benchmark datasets. Our evaluation focuses on the assessment of the generated samples with respect to three important criteria, namely: (1) Machine Learning efficiency, (2) statistical similarity, and (3) privacy risk mitigation. For the task of data imputation, we consider the efficiency of the generated samples across different levels of missing features [^1] Source code will be made available upon acceptance of the manuscript.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/24/05ccs: Data imputation, synthetic data generation, Diffusion Model, Generative Model, Transformerccs: Data imputation, synthetic data generation, Diffusion Model, Generative Model, Transformerccs: Data imputation, synthetic data generation, Diffusion Model, Generative Model, Transformerccs: Data imputation, synthetic data generation, Diffusion Model, Generative Model, Transformer
@@ 1. Introduction
In our analysis, we perform a rigorous evaluation of the proposed approach using several benchmark datasets, each with a wide range of features. We demonstrate that our method consistently outperforms existing baselines, particularly in handling high-dimensional datasets, thereby highlighting its robustness and adaptability in complex tabular data scenarios. The key contributions of this work are the following:
  • We propose a new conditioning mechanism for tabular diffusion models. We model the interaction between condition and masked features (e.g., features to-be-denoised) by using an attention mechanism. Within this mechanism, the embedding of the masked features plays the role of query (Q), and the embedding of the condition plays the roles of key (K) and value (V) (see [54] for details). Compared to the standard approach, where condition and masked features are added [65, 30] or concatenated [32], our method allows to learn more complex relationships showing globally improved performances.
@@
  • We conduct extensive experiments on several public datasets and demonstrate substantial performance gains over state-of-the-art baselines, for both tasks, missing data imputation and synthetic data generation. We evaluate the synthetic data with respect to three important criteria: (1) ML efficiency, (2) statistical similarity and (3) privacy risk.
@@ 2. Related work
Generative models. The application of this family of models to tabular data has been gaining increased attention within the ML community [59, 12, 24, 13, 50, 64, 28, 63, 39, 58]. In particular, tabular VAEs [59] and GANs [59, 12, 24, 13, 50, 64, 28, 63, 39, 58] have shown promising results. Recently, StaSy [29], CoDi [32], Tabsyn [62], TabDDPM [30] have been proposed as powerful alternatives to tabular data generation, leveraging the strengths of Diffusion Models. Specifically, STaSy, applies a score-based generative approach [47], integrating self-paced learning and fine-tuning strategies to enhance data diversity and quality by stabilizing the denoising score matching training process. CoDi addresses training challenges with mixed-data types using a dual diffusion model approach. One model handles continuous features and the other manages discrete (categorical) features. Both models use a UNet-based architecture with linear layers instead of traditional convolutional layers, and are trained to condition on each other’s outputs. This fixed conditioning setup is designed specifically to handle the interactions between continuous and categorical variables effectively. Tabsyn applies the concepts of latent diffusion models [45, 52] to tabular data. First, a VAE encodes mixed-type data into a continuous latent space and then a diffusion model learns this latent distribution. The VAE uses a transformer encoder-decoder to capture feature relationships, while the diffusion model uses an MLP in the reverse denoising process. There are no transformers used in the denoising step, and it does not include a mask conditional method. Finally, TabDDPM manages datasets with mixed data types by using Gaussian diffusion for continuous features and multinomial diffusion for categorical ones. In the preprocessing step, continuous features are scaled using a min-max scaler, and categorical features are one-hot encoded. Then each type of data is sent to its specific diffusion process. After the denoising process, the preprocessing is reversed by scaling back continuous variables, and categorical ones are estimated by first applying a softmax function and then selecting the most likely category. For classification datasets, TabDDPM adopts a class-conditional model consisting in the addition of the condition embedding to the output of the model, while for regression datasets, it integrates target values as an additional feature. It utilizes an MLP architecture as the denosing network optimized with a hybrid objective function that includes mean squared error and KL divergence to predict continuous and categorical data.
@@
Our method is based on TabDDPM, following the same logic of having two separate diffusion models for continuous and categorical data. To this end, we augment the model with three key improvements. First, we employ a transformer-based encoder-decoder as the denoising model, which enhances the capability to learn inter-feature interactions for both, condition and masked features. Second, we integrate the conditioning directly into the transformer’s attention mechanism rather than simply adding embeddings: this approach reduces learning bias, improving how the interaction between the condition and masked features is modeled. Third, we enable dynamic masking during training, which allows our model to handle varying numbers of visible variables, thus supporting both synthetic data generation and missing data imputation within a single framework. In the following sections, we demonstrate that these contributions lead to improved results across various datasets, outperforming TabDDPM and other state-of-the-art algorithms.
@@ 3. Background
The reverse diffusion process $p\left(x_{0:T}\right){=}\prod_{t=1}^{T}p\left(x_{t-1}\middle|x_{t}\right)$ gradually denoises a latent variable $x_{T}{\sim}q\left(x_{T}\right)$ and allows generating new synthetic data. Distributions $p\left(x_{t-1}\middle|x_{t}\right)$ are approximated by a neural network with parameters $\theta$. The parameters are learned optimizing a variational lower bound (VLB):
(1) $L_{\text{vlb}}$ $:=L_{0}+L_{1}+...+L_{T-1}+L_{T}$
(2) $L_{0}$ $:=-\log p_{\theta}(x_{0}|x_{1})$
(3) $L_{T-1}$ $:=D_{KL}(q(x_{T-1}|x_{T},x_{0})\;||\;p_{\theta}(x_{T-1}|x_{T}))$
(4) $L_{T}$ $:=D_{KL}(q(x_{T}|x_{0})\;||\;p(x_{T}))$
The term $q\left(x_{T-1}\middle|x_{T},x_{0}\right)$ is the forward process posterior distribution conditioned on $x_{T}$ and on the initial sample $x_{0}$. $L_{T-1}$ is the Kullback-Leibler divergence between the posterior of the forward process and the parameterized reverse diffusion process $p_{\theta}\left(x_{T-1}\middle|x_{T}\right)$.
Gaussian diffusion models operate in continuous spaces $\left(x_{t}\in\mathbb{R}^{n}\right)$ and in this case the aim of the forward Markov process is to convert the complex unknown data distribution into a known Gaussian distribution. This is achieved by defining a forward noising process $q$ that given a data distribution $x_{0}\sim q(x_{0})$, produces latents $x_{1}$ through $x_{T}$ by adding Gaussian noise at time $t$ with variance $\beta_{t}\in(0,1)$.
@@
(6) $p_{\theta}\left(x_{t-1}\middle|x_{t}\right):=\mathcal{N}\left(x_{t-1};\mu_{\theta}\left(x_{t},t\right),\Sigma_{\theta}\left(x_{t},t\right)\right)$
Ho et al. [18] proposes a simplification of Eq. 6 by employing a diagonal variance $\Sigma_{\theta}\left(x_{t},t\right)=\sigma_{t}I$, where $\sigma_{t}$ are constants dependent on time. This narrows down the prediction task to $\mu_{\theta}\left(x_{t},t\right)$. While a direct prediction of this term via a neural network seems the most intuitive solution, another approach could involve predicting $x_{0}$ and then leveraging earlier equations to determine $\mu_{\theta}\left(x_{t},t\right)$. Alternatively, it could be inferred by predicting the noise $\epsilon$, as done by [18]. In this work, the authors propose the following parameterization:
(7) $\mu_{\theta}\left(x_{t},t\right)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}\left(x_{t},t\right)\right)$
where $\epsilon_{\theta}\left(x_{t},t\right)$ is the prediction of the noise component $\epsilon$ used in the forward diffusion process between the timesteps $t-1$ and $t$, and $\alpha_{t}:=1-\beta_{t},\ \bar{\alpha}_{t}:=\prod_{i\leq t}\alpha_{i}$.
The objective Eq. 1 can be finally simplified to the sum of mean-squared errors between $\epsilon_{\theta}\left(x_{t},t\right)$ and $\epsilon$ over all timesteps $t$:
(8) $L_{\text{simple}}=E_{t,x_{0},\epsilon}\left[||\epsilon-\epsilon_{\theta}(x_{t},t)||^{2}\right]$
@@
Multinomial diffusion models [19] were designed to generate categorical data, where $x_{t}\in\left\{0,1\right\}^{Cl}$ is a one-hot encoded categorical variable with $Cl$ classes. Here, the aim of the forward Markov process is to convert the complex unknown data distribution into a known uniform distribution. The multinomial forward diffusion process $q\left(x_{t}\middle|x_{t-1}\right)$ is a categorical distribution that corrupts the data by uniform noise over $Cl$ classes:
@@
where $\pi=\left[\alpha_{t}x_{t}+(1-\alpha_{t})/Cl\right]\odot\left[\bar{\alpha}_{t-1}x_{0}+(1-\bar{\alpha}_{t-1})/Cl\right]$.
The reverse distribution $p_{\theta}\left(x_{t-1}\middle|x_{t}\right)$ is parameterized as $q\left(x_{t-1}\middle|x_{t},\hat{x}_{0}(x_{t},t)\right)$, where $\hat{x}_{0}$ is predicted by a neural network. Specifically, in this approach, instead of estimating directly the noise component $\epsilon$, we predict $x_{0}$, which is then used to compute the reverse distribution. Then, the model is trained to maximize the VLB Eq. 1.
@@ 4. MTabGen
Although MTabGen shares foundational principles with TabDDPM [30], it differentiates itself through enhancements in both its denoising model and conditioning mechanism. These modifications not only improve synthetic data quality but also enhance the conditioning aspect crucial for the reverse diffusion process. The resulting model excels in both functions: generating conditioned synthetic data and performing missing data imputation.
Figure 1. MTabGen flow for the classification case
@@ 4.1. Problem definition
We focus on tabular datasets for supervised tasks $D=\left\{x^{j_{c}}_{i},x^{j_{n}}_{i},y_{i}\right\}_{i=1}^{N}$ where $x^{j_{n}}_{i}$ with $j_{n}\in\left[1,K_{num}\right]$ is the set of numerical features, $x^{j_{c}}_{i}$ with $j_{c}\in\left[1,K_{cat}\right]$ is the set of categorical features, $y_{i}$ is the label, $i\in\left[1,N\right]$ counts the dataset rows, $N$ is the total number of rows and $K=K_{num}+K_{cat}$ is total number of features.
We apply a consistent preprocessing procedure across our benchmark datasets, using the Gaussian quantile transformation [^2] We have tried different feature normalization methods and obtained similar performance. Using for example Standard Scaler or Min-Max Scaler normalization from scikit-learn, we obtain statistically equivalent results. from the scikit-learn library [41] on numerical features and ordinal encoding for categorical ones. In our approach, we model numerical features with Gaussian diffusion and categorical features with multinomial diffusion. Each feature is subjected to a distinct forward diffusion procedure, which means that the noise components for each feature are sampled individually.
MTabGen generalizes the approach of TabDDPM where the model learns $p\left(x_{t-1}\middle|x_{t},y\right)$, i.e. the probability distribution of $x_{t-1}$ given $x_{t}$ and the target $y$. We extend this by allowing conditioning on a target variable $y$ and a subset of input features, aligning with the strategies proposed by Zheng and Charoenphakdee [65] and Tashiro et al. [49]. Specifically, we partition variable $x$ into $x^{M}$ and $\bar{x}^{M}$. Here, $x^{M}$ refers to the masked variables set, those perturbed by the forward diffusion process, while $\bar{x}^{M}$ represents the untouched variable subset that conditions the reverse diffusion. This setup models $p\left(x^{M}_{t-1}\middle|x^{M}_{t},\bar{x}^{M},y\right)$, with $\bar{x}^{M}$ remaining constant across timesteps $t$. This approach not only enhances model performance in data generation, but it also enables the possibility of performing data imputation with the same model.
The reverse diffusion process $p\left(x^{M}_{t-1}\middle|x^{M}_{t},\bar{x}^{M},y\right)$ is parameterized by the neural network shown in Figs. 1, 2 and 3. In the case of numerical features, the denoising model has to estimate the amount of noise added between steps $t-1$ and $t$ in the forward diffusion process, and in the case of categorical features, it must predict the (logit of) distribution of the categorical variable a $t=0$. The model output has therefore dimensionality of $K_{num}+\sum_{i=1}^{K_{cat}}Cl_{i}$ where $Cl_{i}$ is the number of classes of $i$-th categorical feature.

### v2

@@ Abstract.
Data imputation and data generation have important applications across many domains where incomplete or missing data can hinder accurate analysis and decision-making. Diffusion models have emerged as powerful generative models capable of capturing complex data distributions across various data modalities such as image, audio, and time series. Recently, they have been also adapted to generate tabular data. In this paper, we propose a diffusion model for tabular data that introduces three key enhancements: (1) a conditioning attention mechanism, (2) an encoder-decoder transformer as the denoising network, and (3) dynamic masking. The conditioning attention mechanism is designed to improve the model’s ability to capture the relationship between the condition and synthetic data. The transformer layers help model interactions within the condition (encoder) or synthetic data (decoder), while dynamic masking enables our model to efficiently handle both missing data imputation and synthetic data generation tasks within a unified framework. We conduct a comprehensive evaluation by comparing the performance of diffusion models with transformer conditioning against state-of-the-art techniques such as Variational Autoencoders, Generative Adversarial Networks and Diffusion Models, on benchmark datasets. Our evaluation focuses on the assessment of the generated samples with respect to three important criteria, namely: (1) Machine Learning efficiency, (2) statistical similarity, and (3) privacy risk mitigation. For the task of data imputation, we consider the efficiency of the generated samples across different levels of missing features. The results demonstrates average superior machine learning efficiency and statistical accuracy compared to the baselines, while maintaining privacy risks at a comparable level, particularly showing increased performance in datasets with a large number of features. By conditioning the data generation on a desired target variable, the model can mitigate systemic biases, generate augmented datasets to address data imbalance issues, and improve data quality for subsequent analysis. This has significant implications for domains such as healthcare and finance, where accurate, unbiased, and privacy-preserving data are critical for informed decision-making and fair model outcomes  [^1] Source code will be made available upon acceptance of the manuscript.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/24/05ccs: Computer systems organization Neural networksccs: Computing methodologies Temporal difference learningccs: Computing methodologies Learning latent representationsccs: Computing methodologies Machine learning approachesccs: Computing methodologies Machine learningccs: Computing methodologiesccs: Computing methodologies Neural networks
@@ 1. Introduction
In our analysis, we perform a rigorous evaluation of the proposed approach using several benchmark datasets, each with a wide range of features. We demonstrate that our method shows overall improved performance compared to existing baselines, particularly in handling high-dimensional datasets, thereby highlighting its robustness and adaptability in complex tabular data scenarios. The key contributions of this work are the following:
  • We propose a new conditioning mechanism for tabular diffusion models. We model the interaction between condition and masked features (e.g., features to-be-denoised) by using an attention mechanism. Within this mechanism, the embedding of the masked features plays the role of query ($\bf{Q}$), and the embedding of the condition plays the roles of key ($\bf{K}$) and value ($\bf{V}$) (see [59] for details). Compared to the standard approach, where condition and masked features are added [70, 32] or concatenated [34], our method allows to learn more complex relationships showing globally improved performances.
@@
  • We conduct extensive experiments on several public datasets and demonstrate average performance gain over state-of-the-art baselines, for both tasks, missing data imputation and synthetic data generation. We evaluate the synthetic data with respect to three important criteria: (1) ML efficiency, (2) statistical similarity and (3) privacy risk.
@@ 2. Related work
Tabular Data Modeling and Benchmarks. The evaluation of models for tabular data requires diverse benchmarks that reflect real-world data conditions. Several recent studies have established such benchmarks for both data imputation and supervised learning. Jäger et al. [23] present a comprehensive evaluation of data imputation methods, comparing both classical and deep learning-based imputation approaches. Their results show that imputation can significantly improve downstream ML tasks, especially when training data is fully observed. For supervised tasks, Grinsztajn et al. [16] argue that tree-based models such as XGBoost and Random Forests often outperform deep learning models due to the specific properties of tabular data, including irregular patterns and uninformative features. The authors provide a standard benchmark for evaluating both traditional and novel approaches on a variety of real-world datasets. Borisov et al. [8] also highlight the challenges faced by deep learning models when dealing with heterogeneous tabular data and emphasize the need for architectures specifically designed for tabular datasets. In this work, we extend the benchmark proposed by [8] to include a wider variety of datasets, providing a more exhaustive evaluation of generative tasks.
@@
Generative models. The application of this family of models to tabular data has been gaining increased attention within the ML community. In particular, tabular VAEs [64] and GANs [13, 55, 69, 30, 68, 42, 63] have shown promising results. Recently, StaSy [31], CoDi [34], Tabsyn [67], TabDDPM [32] have been proposed as powerful alternatives to tabular data generation, leveraging the strengths of Diffusion Models.
Specifically, STaSy, applies a score-based generative approach [52], integrating self-paced learning and fine-tuning strategies to enhance data diversity and quality by stabilizing the denoising score matching training process.
CoDi addresses training challenges with mixed-data types using a dual diffusion model approach. One model handles continuous features and the other manages discrete (categorical) features. Both models use a UNet-based architecture with linear layers instead of traditional convolutional layers, and are trained to condition on each other’s outputs. This fixed conditioning setup is designed specifically to handle the interactions between continuous and categorical variables effectively.
Tabsyn applies the concepts of latent diffusion models [49, 57] to tabular data. First, a VAE encodes mixed-type data into a continuous latent space and then a diffusion model learns this latent distribution. The VAE uses a transformer encoder-decoder to capture feature relationships, while the diffusion model uses an MLP in the reverse denoising process. There are no transformers used in the denoising step, and it does not include a mask conditional method.
@@
Finally, TabDDPM manages datasets with mixed data types by using Gaussian diffusion for continuous features and multinomial diffusion for categorical ones. In the preprocessing step, continuous features are scaled using a min-max scaler, and categorical features are one-hot encoded. Then each type of data is sent to its specific diffusion process. After the denoising process, the preprocessing is reversed by scaling back continuous variables, and categorical ones are estimated by first applying a softmax function and then selecting the most likely category. For classification datasets, TabDDPM adopts a class-conditional model consisting in the addition of the condition embedding to the output of the model, while for regression datasets, it integrates target values as an additional feature. It utilizes an MLP architecture as the denosing network optimized with a hybrid objective function that includes mean squared error and KL divergence to predict continuous and categorical data.
@@
Our method is based on TabDDPM, following the same logic of having two separate diffusion models for continuous and categorical data. To this end, we augment the model with three key improvements. First, we employ a transformer-based encoder-decoder as the denoising model, which enhances the capability to learn inter-feature interactions for both, condition and masked features. Second, we integrate the conditioning directly into the transformer’s attention mechanism rather than simply adding embeddings: this approach reduces learning bias, improving how the interaction between the condition and masked features is modeled. Third, we enable dynamic masking during training, which allows our model to handle varying numbers of visible variables, thus supporting both synthetic data generation and missing data imputation within a single framework. In the following sections, we demonstrate that these contributions lead to improved results across various datasets, generally outperforming TabDDPM and other state-of-the-art algorithms.
@@ 3. Background
The reverse diffusion process $p\left(x_{0:T}\right){=}\prod_{t=1}^{T}p\left(x_{t-1}\middle|x_{t}\right)$ gradually denoises a latent variable $x_{T}{\sim}q\left(x_{T}\right)$ and allows generating new synthetic data. Distributions $p\left(x_{t-1}\middle|x_{t}\right)$ are approximated by a neural network with parameters $\theta$.
In this work, we use the hat notation (e.g., $\hat{p}\left(x_{t-1}\middle|x_{t}\right)$) to indicate that a variable is estimated by a neural network model trained with parameters $\theta$. Although these estimations depend on $\theta$, we omit $\theta$ for notational simplicity. Thus, $\hat{p}\left(x_{t-1}\middle|x_{t}\right)$ should be interpreted as $\hat{p}_{\theta}\left(x_{t-1}\middle|x_{t}\right)$, where the model parameters $\theta$ influence the predicted value.
The parameters are learned optimizing a variational lower bound (VLB):
(1) $L_{\text{vlb}}$ $:=\sum_{t=0}^{T}L_{t}$
(2) $L_{0}$ $:=-\log\hat{p}\left(x_{0}\middle|x_{1}\right)$
(3) $L_{t-1}$ $:=D_{KL}(q\left(x_{t-1}\middle|x_{t},x_{0}\right)\;||\;\hat{p}\left(x_{t-1}\middle|x_{t}\right))$
(4) $L_{T}$ $:=D_{KL}(q\left(x_{T}\middle|x_{0}\right)\;||\;p\left(x_{T}\right))$
The term $q\left(x_{t-1}\middle|x_{t},x_{0}\right)$ is the forward process posterior distribution conditioned on $x_{t}$ and on the initial sample $x_{0}$. $L_{t-1}$ is the Kullback-Leibler divergence between the posterior of the forward process and the estimated reverse diffusion process $\hat{p}\left(x_{t-1}\middle|x_{t}\right)$.
Gaussian diffusion models operate in continuous spaces $\left(x_{t}\in\mathbb{R}^{n}\right)$ and in this case the aim of the forward Markov process is to convert the complex unknown data distribution into a known Gaussian distribution. This is achieved by defining a forward noising process $q$ that given a data distribution $x_{0}\sim q\left(x_{0}\right)$, produces latents $x_{1}$ through $x_{T}$ by adding Gaussian noise at time $t$ with variance $\beta_{t}\in(0,1)$.
@@
(6) $\hat{p}\left(x_{t-1}\middle|x_{t}\right):=\mathcal{N}\left(x_{t-1};\hat{\mu}\left(x_{t},t\right),\hat{\Sigma}\left(x_{t},t\right)\right)$
Ho et al. [19] proposes a simplification of Eq. 6 by employing a diagonal variance $\hat{\Sigma}\left(x_{t},t\right)=\sigma_{t}I$, where $\sigma_{t}$ are constants dependent on time. This narrows down the prediction task to $\hat{\mu}\left(x_{t},t\right)$. While a direct prediction of this term via a neural network seems the most intuitive solution, another approach could involve predicting $x_{0}$ and then leveraging earlier equations to determine $\hat{\mu}\left(x_{t},t\right)$. Alternatively, it could be inferred by predicting the noise $\epsilon$, as done by Ho et al. [19]. In this work, the authors propose the following parameterization:
(7) $\hat{\mu}\left(x_{t},t\right)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon}\left(x_{t},t\right)\right)$
where $\hat{\epsilon}\left(x_{t},t\right)$ is the prediction of the noise component $\epsilon$ used in the forward diffusion process between the timesteps $t-1$ and $t$, and $\alpha_{t}:=1-\beta_{t},\ \bar{\alpha}_{t}:=\prod_{i\leq t}\alpha_{i}$.
The objective Eq. 1 can be finally simplified to the sum of mean-squared errors between $\hat{\epsilon}\left(x_{t},t\right)$ and $\epsilon$ over all timesteps $t$:
(8) $L_{\text{simple}}=E_{t,x_{0},\epsilon}\left[\left|\left|\epsilon-\hat{\epsilon}\left(x_{t},t\right)\right|\right|^{2}\right]$
@@
Multinomial diffusion models Hoogeboom et al. [20] designed the procedures for generating categorical data, where $x_{t}\in\left\{0,1\right\}^{Cl}$ is a one-hot encoded categorical variable with $Cl$ classes. Here, the aim of the forward Markov process is to convert the complex unknown data distribution into a known uniform distribution. The multinomial forward diffusion process $q\left(x_{t}\middle|x_{t-1}\right)$ is a categorical distribution that corrupts the data by uniform noise over $Cl$ classes:
@@
where $\pi=\left[\alpha_{t}x_{t}+\left(1-\alpha_{t}\right)/Cl\right]\odot\left[\bar{\alpha}_{t-1}x_{0}+\left(1-\bar{\alpha}_{t-1}\right)/Cl\right]$.
The reverse distribution $\hat{p}\left(x_{t-1}\middle|x_{t}\right)$ is parameterized as $q\left(x_{t-1}\middle|x_{t},\hat{x}_{0}(x_{t},t)\right)$, where $\hat{x}_{0}$ is predicted by a neural network. Specifically, in this approach, instead of estimating directly the noise component $\epsilon$, we predict $x_{0}$, which is then used to compute the reverse distribution. Then, the model is trained to maximize the VLB Eq. 1.
@@ 4. MTabGen
MTabGen shares foundational principles with the TabDDPM approach [32], but introduces enhancements in both the denoising model and the conditioning mechanism using a transformer encoder-decoder architecture. These improvements enhance the quality of synthetic data and strengthen the conditioning process needed for the reverse diffusion. As a result, the model is capable of generating conditioned synthetic data and performing missing data imputation, generally improving on the performance of other state-of-the-art models for these tasks.
@@ 4.1. Problem definition
In this work, we address the challenge of modeling tabular datasets for supervised tasks. Our approach is designed to handle both numerical and categorical features, while also effectively dealing with missing values. We focus on tabular datasets for supervised tasks $D=\left\{\left(x^{c}_{i},x^{n}_{i},y_{i}\right)\right\}_{i=1}^{N}$, where $x^{n}_{i}\in\mathbb{R}^{K_{\text{num}}}$ represents the set of numerical features, $x^{c}_{i}\in\mathbb{Z}^{K_{\text{cat}}}$ represents the set of categorical features, with each categorical feature potentially having a different number of categories, and $y_{i}$ is the target label for row $i$. The dataset contains $N$ rows, where $K_{\text{num}}$ and $K_{\text{cat}}$ are the number of numerical and categorical features respectively, with the total number of features being $K=K_{\text{num}}+K_{\text{cat}}$.
In our approach, we model numerical features with Gaussian diffusion and categorical features with multinomial diffusion. Each feature is subjected to a distinct forward diffusion procedure, which means that the noise components for each feature are sampled individually.
MTabGen generalizes the approach of TabDDPM where the model originally learned $p\left(x_{t-1}\middle|x_{t},y\right)$, i.e. the probability distribution of $x_{t-1}$ given $x_{t}$ and the target $y$. We extend this by allowing conditioning not only on the target $y$ but also on a subset of input features, aligning with the strategies proposed by Zheng and Charoenphakdee [70] and Tashiro et al. [54]. Specifically, we partition variable $x$ into two subsets: $x^{\text{mask}}$ and $x^{\text{cond}}$. Here, $x^{\text{mask}}$ contains the features masked and subjected to forward diffusion, while $x^{\text{cond}}$ represents the untouched variable subset that conditions the reverse diffusion. This setup models $p\left(x^{\text{mask}}_{t-1}\middle|x^{\text{mask}}_{t},x^{\text{cond}},y\right)$, with $x^{\text{cond}}$ and $y$ remaining constant across timesteps $t$. This approach not only enhances model performance in data generation, but it also enables the possibility of performing data imputation with the same model.
The reverse diffusion process $p\left(x^{\text{mask}}_{t-1}\middle|x^{\text{mask}}_{t},x^{\text{cond}},y\right)$ is parameterized by a single neural network shown in Figs. 1, 2 and 3. This common neural network predicts simultaneously both the amount of noise added for numerical features between steps $t-1$ and $t$ and the distribution of categorical features at time $t=0$. The output dimensionality is $K_{\text{num}}+\sum_{i=1}^{K_{\text{cat}}}Cl_{i}$, where $Cl_{i}$ denotes the number of classes for the $i$-th categorical feature.